determine weather the following mappings are linear transformations. Either prove that the mapping is a linear transformation to explain why it is not a linear transformation.
a)T:R3[x] to R3[x] given by T(p(x))=xp'(x)+1, where f'(x) is a derivative of the polynomial p(x).
b) T:R2 to R2 given by T([x y])=[x -y]. Additionally describe this mapping in part b geometrically.
Answer #1

در |=( F ( x x ) = * )ط + (a2 اطلا= ((عط)T ر+ () لا = 20 +1 = (28 ) - ((2px) آ F | (2) 3 = (203 الجدد عطر (ر) ط 2)T lineen. t